Find the center (h,k)and the radius of the circle using this equation: 5x^2+8x+5y^2+5y-9=0
Answer by lwsshak3(11628)

Find the center (h,k)and the radius of the circle using this equation:
5x^2+8x+5y^2+5y-9=0
**
equation of circle: (x-h)^+(y-k)^2=r^2, (h,k)=(x,y) coordinates of center, r=radius
5x^2+8x+5y^2+5y-9=0
complete the squares
5(x^2+8x/5+64/100)+5(y^2+y+(1/4))=9+64/20+5/4
5(x+(8/10)^2)+5(y+(1/2)^2=180/20+64/20+25/20=269/20
5(x+(8/10)^2)+5(y+(1/2)^2=269/20
divide both sides by 5
(x+(8/10)^2)+(y+(1/2)^2=269/100
..
center: (-4/5,-1/2)
radius:√(269/100)=√269/10≈1.64